# Short-term plasticity recalls forgotten memories through a trampoline mechanism

### v1

@@ IV.2 Stable fixed points from short-term synaptic plasticity
The four blocks $\mathcal{J}_{xx},\,\mathcal{J}_{xA},\,\mathcal{J}_{Ax},\,\mathcal{J}_{AA}$ correspond to neuron-neuron, neuron-synapses, synapses-neuron, and synapses-synapses perturbations. Following Clark and Abbott [clark2024dynamic], we study the salient aspects of the spectrum of the Jacobian $\mathcal{J}$ (details in App. F). The spectrum of $\mathcal{J}$ has $N^{2}-N$ degenerate eigenvalues $\lambda=-1/p$, which correspond to perturbations belonging entirely to the space of the short-term, plastic couplings $A_{ij}$. The remaining $2N$ eigenvalues are determined by the solutions of the equation
@@
Here, $\mathbf{I}_{N}$ is the $N\times N$ identity matrix, $\bm{\phi}^{*}$ is a $N$-dimensional vector containing the output of the neurons at a fixed point of the neural network dynamics, and $\Phi^{\prime}]_{ij}={\phi^{*}_{i}}^{\prime}\delta_{ij}$ is a diagonal matrix whose diagonal elements are the derivatives of the output of the neurons at the fixed point soft the dynamics. The quantity $q_{\infty}=\frac{1}{N}\sum_{i=1}^{N}\langle(\phi^{*}_{i})^{2}\rangle$ is the self-overlap of the neuron output at the fixed point, and it is equivalent to the long time limit $\lim_{t\to\infty}C_{\phi}(t,t)$ of the autocorrelation function in Eq. (21) from the DMFT. Finally, the matrix $\mathbf{W}^{*}$ contains the long term and short-term synaptic weights at the fixed point, $\mathbf{W}^{*}]_{ij}=J_{ij}+A^{*}_{ij}$.
The resulting eigenvalue structure, namely the $2N$ coupled neuron–synapse eigenvalues obtained from Eq. (24) together with the $N^{2}-N$ degenerate eigenvalues $\lambda=-1/p$ coming from $\mathcal{J}_{AA}$, clarifies how short-term synaptic plasticity affects the nature of the excitations around fixed points of the landscape. In the extended system composed by neurons and synapses, the block $\mathcal{J}_{AA}$ always contributes $N^{2}$ eigenvalues at 1/p-1/p, corresponding to the intrinsic relaxation of the synapses. When $k=0$, the coupling matrix $\mathbf{C}$ vanishes, so these synaptic modes are completely decoupled from the neural dynamics, and the stability of the neuronal fixed point is determined solely by the spectrum of $\mathcal{J}_{xx}$. For $k>0$, however, the plasticity-induced correction $\mathbf{C}$ provides an additional feedback channel through which activity perturbations influence the neural dynamics. This extra dissipative pathway mirrors the intuition from the analysis of the dynamics in App. C: synaptic dynamics suppress or soften unstable modes that would otherwise destabilize the purely static network. Finally, in the limit of very large $k$, the fixed points typically lie in a saturated regime of the neuronal nonlinearity, so that $\phi^{\prime}(x_{i}^{*})\approx 0$. In this case $\Phi^{\prime}\to\mathbf{0}$ and the effective feedback matrices $\mathcal{J}_{xx}=-\mathbf{I}_{N}+\mathbf{W}^{*}\Phi^{\prime}$ and $\mathbf{C}\propto\Phi^{\prime}$ reduce to their leak terms, yielding a spectrum of eigenvalues concentrated at 1-1 (neurons) and 1/p-1/p (synapses). Thus, linear excitations affecting neurons and synapses become effectively decoupled.
@@ IV.3 Optimal synaptic timescale
In this Section, we discuss the effects of the short-term plasticity timescale $p$ of the plastic dynamics on the retrieval capability of the network. Indeed, from our DMFT analysis in Sec. IV.1, we see that short-term synaptic plasticity affects the dynamics by means of a feedback mechanism, whose signals decays over a timescale $p$. If the plastic dynamics is too fast, the synapses $A_{ij}$ track the neural activity too closely, preventing the formation of a stable representation. If the plastic dynamics is too slow, the plastic synapses cannot react in time to stabilize the transient neural dynamics, and catastrophic forgetting is expected to occur. This qualitative analysis suggests the existence of an optimal timescale over which short-term synaptic plasticity should act in order to robustly enlarge the basin of attraction around a desired memory. This is precisely what we demonstrate in Fig. 5.
@@
For a fixed load $\alpha>\alpha_{c}$ we solve numerically the DMFT equations Eqs. 16, 17, 18, 19, 20, 21 and 22 for different values of $p$, and determine the asymptotic overlap of the network with the target pattern $m^{*}_{\infty}$. We then plot the asymptotic value of the overlap, as a function of the short-term plasticity timescale $p$. Different curves are constructed for different values of the short-term plasticity strength $k$. They all display a nonmonotonic behavior as the characteristic short-term plasticity timescale $p$ is varied. All the curves reach a maximum around $p\approx 2$, with the location of the maximum exhibiting only a weak dependence on $k$. A comparison with Fig. 3 (a) shows that the plastic timescale that allows optimal plastic retrieval when $k>0$ roughly matches the time at which the maximum overlap is achieved at $k=0$ during the transient recovery dynamics. This finding corroborates the picture that short-term synaptic plasticity exploits the transient recovery trajectories to lead the neural network toward stable fixed points in the vicinity of the target pattern. Interestingly, we observe that the plastic timescale does not enter at all into the expression of the Lyapunov function $\mathcal{L}$ in Eq. (7), which describes the energy landscape in which the network evolve: the metastable states created by short-term synaptic plasticity have thus a complex structure, and the ability of the network to fall into them can be enhanced or suppressed by changing the short-term plasticity timescale.
@@ V Outlook
In this work, we studied how short-term associative synaptic plasticity affects the retrieval capabilities of a recurrent neural network with a long-term Hebbian learning rule. Short-term synaptic plasticity propagates past neural activity to the current state of the network, and it constantly reshapes the effective landscape in which the neurons evolve. The evolution of the neurons in this restricted landscape is thus akin to the motion of a ball on an elastic trampoline, constantly reshaped by the ball’s position. By this means, the network is able to recover memories that would be otherwise forgotten, and the network’s retrieval abilities are substantially enhanced. We conclude by outlining interesting directions our work opens the way to.
@@
First, inspecting the impact of short-term synaptic plasticity on retrieval properties of dense associative memories [krotov2016dense, lucibello2024exponential] is a relevant question with applications in both biologically plausible learning and artificial neural network. Due to the exponentially large capacity of dense associative memories, any improvement in pattern retrieval capabilities can yield a great impact on memory storage. It is unclear at the moment, and worth investigating, whether the computational workload implied by the update of dense synapses can be circumvented simply by the addition of sparser realizations of short-term synaptic plasticity.
@@
Second, another intriguing direction consists in achieving a fine-grained description of the impact of short-term synaptic plasticity on the basins of attraction of recurrent neural networks. The fact that the dynamics of the system optimize an energy function allows the deployment of techniques from the statistical mechanics of spin-glasses to characterize these metastable states, such as the Franz-Parisi potential [franz1995recipes] or the Kac-Rice method [fyodorov2004complexity, ros2019complex]. However, the complex spin-glass structure of the recurrent neural network poses impressive technical challenges, already without short-term synaptic plasticity, but achieving such an understanding is a very much desirable long-term objective.
@@
Third, it would be interesting to understand the interplay between short-term synaptic plasticity and noise in the neural network. In biological models this noise could stem from neural activity in nearby regions, and recent work has shown that adding an out of equilibrium noise can increase significantly the storage capacity of the network [behera2023enhanced, du2024active]. How does noise affects short-term synaptic plasticity, and how do the qualitative features of this effect change when the ratio between the characteristic timescale of noise and short-term synaptic plasticity are varied is an interesting question to address.
@@
Finally, a direction worth investigating consists in understanding how can short-term synaptic plasticity reshape the landscape of Hopfield-like recurrent neural network if an external drive is imposed to the system and both short-term plasticity and short-term antiplasticity (the latter can be realized using a negative $k$) are alternatively allowed, before freezing the whole set of synaptic weights of the network, in the spirit of what has been done in [wakhloo2025associative]. Is it possible to erase one of the long-term stored patterns, and replace it with a new one? If a dynamical stimulus is applied to our recurrent neural network, can limit cycle and complex trajectories be stored within the basin of attraction of a target pattern, or would they destroy the metastable states generated through short-term synaptic plasticity? Progress on these questions should be possible building on the aforementioned study and our work.
@@ Acknowledgements.
We thank Itamar D. Landau, Atshushi Yamamura and Aditya Mahadevan for insightful discussions. F.G. acknowledges support from the Stanford Leinweber’s Institute of Theoretical Physics. S.G. thanks the Simons foundation and a Schmidt foundation polymath award for support. We thank the Stanford Sherlock cluster for computational resources.
Appendix A Lyapunov function of RNN with short-term synaptic plasticity
@@
We firstly differentiate the WW–dependent term. We obtain
@@
Appendix B Cavity Calculation of Global Minima
In this Appendix, we use the static cavity method [mezard1987, del2014cavity] to obtain a set of self-consistent equations for the fixed point of the recurrent neural network with short-term synaptic plasticity and long-term Hebbian rule. Our starting point are the equation of motion for the action potential of the neurons in the system and the plasticity matrix, given by Eq. (1) and Eq. (5). These equations are recapitulated here for completeness, and are given by
@@ Appendix B Cavity Calculation of Global Minima
B.1 Breakdown of the cavity and spin-glass transition
@@
Expression for the mean squared susceptibilities $\overline{\chi}^{(m,s)}_{2}$, $\overline{\chi}^{(\phi,h)}_{2}$ can be derived through a similar procedure to the one followed in Sec. B for the derivation of the susceptibilities. The final result is
@@
B.2 Phase boundaries with short-term synaptic plasticity
@@
We now address how short-term synaptic plasticity affects the shape of the boundary $\gamma_{g}(\alpha)$ between the paramagnetic phase and the spin glass phase. This transition corresponds to the point where the single-site cavity description breaks down due to the divergence of the fragility $\Xi$, as discussed in Appendix B.1. The condition for this breakdown is given by Eq. (64), when the inequality is saturated. In the paramagnetic phase ($m^{*}=q=0$), the membrane potential of a cavity neuron is $x=\Gamma\phi$ (see Eqs. 9, 10, 10, 11, 12, 13 and 14). Since the membrane potential is independent from the realizations of the Gaussian field $z$, we have that $\left[\overline{\chi}^{(\phi,h)}\right]^{2}=\overline{\chi}_{2}^{(\phi,h)}$. Therefore, the instability condition in Eq. (64) becomes
@@
Appendix C DMFT Derivation
@@
since $A_{ij}(0)=0$, only the non-local self-interaction term survives, similarly to the results obtained by [clark2024dynamic]. Inserting Eq. (77) into Eq. (1), we obtain the following expression:
@@
Since we already dealt with the synaptic term, this can be left out in the following computations. We will focus on the term with $J_{ij}$ only, and put back the synaptic kernel in the end. As done in App. B, we introduce a dynamical overlap for each pattern $\mu=1,\ldots,P$,
@@
and we split the set of patterns $\xi_{i}^{\mu}$ into a set of condensed patterns, denoted without loss of generality by $\{\mu^{\ast}\}=\{1,\dots,s\}$, and a set of uncondensed patterns, for $\mu>s$. The overlap against condensed patterns are of order $m^{\mu^{*}}(t)\sim O(1)$, while the overlaps against the uncondensed patterns are of order $m^{\mu}(t)\sim O(N^{-1/2})$. We are going to add in two distinct steps two distinct cavity variables (a neuron and a pattern respectively, as done in App. B) to the dynamical evolution of a neuron in the absence of short-term synaptic plasticity:
@@
C.1 Final self-consistent DMFT with plastic kernel

### v2

@@ IV.2 Stable fixed points from short-term synaptic plasticity
The four blocks $\mathcal{J}_{xx},\,\mathcal{J}_{xA},\,\mathcal{J}_{Ax},\,\mathcal{J}_{AA}$ correspond to neuron-neuron, neuron-synapses, synapses-neuron, and synapses-synapses perturbations. Following Clark and Abbott [7], we study the salient aspects of the spectrum of the Jacobian $\mathcal{J}$ (details in App. G). The spectrum of $\mathcal{J}$ has $N^{2}-N$ degenerate eigenvalues $\lambda=-1/p$, which correspond to perturbations belonging entirely to the space of the short-term, plastic couplings $A_{ij}$. The remaining $2N$ eigenvalues are determined by the solutions of the equation
@@
Here, $\mathbf{I}_{N}$ is the $N\times N$ identity matrix, $\bm{\phi}^{*}$ is a $N$-dimensional vector containing the output of the neurons at a fixed point of the neural network dynamics, and $\Phi^{\prime}]_{ij}={\phi^{*}_{i}}^{\prime}\delta_{ij}$ is a diagonal matrix whose diagonal elements are the derivatives of the output of the neurons at the fixed point of the dynamics. The quantity $q_{\infty}=\frac{1}{N}\sum_{i=1}^{N}\langle(\phi^{*}_{i})^{2}\rangle$ is the self-overlap of the neuron output at the fixed point, and it is equivalent to the long time limit $\lim_{t\to\infty}C_{\phi}(t,t)$ of the autocorrelation function in Eq. (21) from the DMFT. Finally, the matrix $\mathbf{W}^{*}$ contains the long term and short-term synaptic weights at the fixed point, $\mathbf{W}^{*}]_{ij}=J_{ij}+A^{*}_{ij}$.
The resulting eigenvalue structure, namely the $2N$ coupled neuron–synapse eigenvalues obtained from Eq. (24) together with the $N^{2}-N$ degenerate eigenvalues $\lambda=-1/p$ coming from $\mathcal{J}_{AA}$, clarifies how short-term synaptic plasticity affects the nature of excitations around fixed points of the landscape. In the extended system composed by neurons and synapses, the block $\mathcal{J}_{AA}$ always contributes $N^{2}$ eigenvalues at 1/p-1/p, corresponding to the intrinsic relaxation of the synapses. When $k=0$, the coupling matrix $\mathbf{C}$ vanishes, so these synaptic modes are completely decoupled from the neural dynamics, and the stability of the neuronal fixed point is determined solely by the spectrum of $\mathcal{J}_{xx}$. For $k>0$, however, the plasticity-induced correction $\mathbf{C}$ provides an additional feedback channel through which activity perturbations influence the neural dynamics. This extra dissipative pathway mirrors the intuition from the analysis of the dynamics in App. D: synaptic dynamics suppress or soften unstable modes that would otherwise destabilize the purely static network. Finally, in the limit of very large $k$, the fixed points typically lie in a saturated regime of the neuronal nonlinearity, so that $\phi^{\prime}(x_{i}^{*})\approx 0$. In this case $\Phi^{\prime}\to\mathbf{0}$ and the effective feedback matrices $\mathcal{J}_{xx}=-\mathbf{I}_{N}+\mathbf{W}^{*}\Phi^{\prime}$ and $\mathbf{C}\propto\Phi^{\prime}$ reduce to their leak terms, yielding a spectrum of eigenvalues concentrated at 1-1 (neurons) and 1/p-1/p (synapses). Thus, linear excitations affecting neurons and synapses become effectively decoupled.
@@ IV.3 Optimal synaptic timescale
In this Section, we discuss the effects of the short-term plasticity timescale $p$ of the plastic dynamics on the retrieval capability of the network. Indeed, from our DMFT analysis in Sec. IV.1, we see that short-term synaptic plasticity affects the dynamics by means of a feedback mechanism, whose signal decays over a timescale $p$. If the plastic dynamics is too fast, the synapses $A_{ij}$ track the neural activity too closely, preventing the formation of a stable representation. If the plastic dynamics is too slow, the plastic synapses cannot react in time to stabilize the transient neural dynamics, and catastrophic forgetting is expected to occur. This qualitative analysis suggests the existence of an optimal timescale over which short-term synaptic plasticity should act in order to robustly enlarge the basin of attraction around a desired memory. This is precisely what we demonstrate in Fig. 5.
@@
For a fixed load $\alpha>\alpha_{c}$ we solve numerically the DMFT equations Eqs. 16, 17, 18, 19, 20, 21 and 22 for different values of $p$, and determine the asymptotic overlap of the network with the target pattern $m^{*}_{\infty}$. We then plot the asymptotic value of the overlap, as a function of the short-term plasticity timescale $p$. Different curves are constructed for different values of the short-term plasticity strength $k$. They all display a nonmonotonic behavior as the characteristic short-term plasticity timescale $p$ is varied. All the curves reach a maximum around $p\approx 2$, with the location of the maximum exhibiting only a weak dependence on $k$. A comparison with Fig. 3 (a) shows that the plastic timescale that allows optimal plastic retrieval when $k>0$ roughly matches the time at which the maximum overlap is achieved at $k=0$ during the transient recovery dynamics. This finding corroborates the picture that short-term synaptic plasticity exploits the transient recovery trajectories to lead the neural network toward stable fixed points in the vicinity of the target pattern. Interestingly, we observe that the plastic timescale does not enter at all into the expression of the Lyapunov function $\mathcal{L}$ in Eq. (7), which describes the energy landscape in which the network evolves: the metastable states created by short-term synaptic plasticity have thus a complex structure, and the ability of the network to fall into them can be enhanced or suppressed by changing the short-term plasticity timescale.
@@ V Outlook
In this work, we studied how short-term associative synaptic plasticity affects the retrieval capabilities of a recurrent neural network with a long-term Hebbian learning rule. Short-term synaptic plasticity propagates past neural activity to the current state of the network, and it constantly reshapes the effective landscape in which the neurons evolve. The evolution of the neurons in this restricted landscape is thus akin to the motion of a ball on an elastic trampoline, constantly reshaped by the ball’s position. By this means, the network is able to recover memories that would be otherwise forgotten, and the network’s retrieval abilities are substantially enhanced. We conclude by outlining interesting directions that our work opens the way to.
@@
First, inspecting the impact of short-term synaptic plasticity on retrieval properties of dense associative memories [28, 34] is a relevant question with applications in both biologically plausible learning and artificial neural network. Due to the exponentially large capacity of dense associative memories, any improvement in pattern retrieval capabilities can yield a great impact on memory storage. It is unclear at the moment, and worth investigating, whether the computational workload implied by the update of dense synapses can be circumvented simply by the addition of sparser realizations of short-term synaptic plasticity.
@@
Second, another intriguing direction consists in achieving a fine-grained description of the impact of short-term synaptic plasticity on the basins of attraction of recurrent neural networks. The fact that the dynamics of the system optimize an energy function allows the deployment of techniques from the statistical mechanics of spin-glasses to characterize these metastable states, such as the Franz-Parisi potential [15] or the Kac-Rice method [16, 46]. However, the complex spin-glass structure of the recurrent neural network poses impressive technical challenges, already without short-term synaptic plasticity, but achieving such an understanding is a very much desirable long-term objective.
@@
Third, it would be interesting to understand the interplay between short-term synaptic plasticity and noise in the neural network. In biological models this noise could stem from neural activity in nearby regions, and recent work has shown that adding an out of equilibrium noise can increase significantly the storage capacity of the network [6, 11]. How does noise affects short-term synaptic plasticity, and how do the qualitative features of this effect change when the ratio between the characteristic timescale of noise and short-term synaptic plasticity are varied is an interesting question to address.
@@
Finally, a direction worth investigating consists in understanding how short-term synaptic plasticity reshapes the landscape of Hopfield-like recurrent neural network if an external drive is imposed to the system and both short-term plasticity and short-term antiplasticity (the latter can be realized using a negative $k$) are alternatively allowed, before freezing the whole set of synaptic weights of the network, in the spirit of what has been done in [55]. Is it possible to erase one of the long-term stored patterns, and replace it with a new one? If a dynamical stimulus is applied to the network, can limit cycles and complex trajectories be stored within the basin of attraction of a target pattern, or would they destroy the metastable states generated through short-term synaptic plasticity? Progress on these questions should be possible on the basis of the aforementioned study and our work.
@@ Acknowledgements.
We thank Itamar D. Landau, Aditya Mahadevan and Atshushi Yamamura for insightful discussions. F.G. acknowledges support from the Stanford Leinweber’s Institute of Theoretical Physics. S.G. thanks the Simons foundation and a Schmidt foundation polymath award for support. We thank the Stanford Sherlock cluster for computational resources.
Appendix A Details on Fig. 1(b–d) and Fig. 1(f–h)
In this Appendix, we report the numerical parameters and implementation details used to generate the energy–landscape plots shown in Fig. 1 (b–d) and Fig. 1 (f–h), where the dynamics of a low-dimensional neural network is simulated, in a parameter regime that allow us to mimic the effects of transient retrieval, catastrophic forgetting, and plastic retrieval, while illustrating the trampoline mechanism induced by short-term synaptic plasticity.
The system consists of $N=2$ continuous neurons, with two stored binary patterns: a ”spurious” pattern $\bm{\xi}^{(1)}=(1,-1)$ and a ”target” pattern $\bm{\xi}^{(2)}=(-1,1)$. The position of the latter is identified by the red cross in Fig. 1. The fixed Hopfield coupling matrix $\mathbf{J}$ is constructed from these patterns as described in Eq. (4). The system follows the dynamics given by Eq. (1) and Eq. (5), with the addition of an external field $h\xi_{i}^{(1)}$ on the right hand side of Eq. (1). The field points along the direction of the spurious pattern, and it is included to destabilize the target memory. Because of this additional magnetic field, the Lyapunov function $\mathcal{L}$ of the 2-neurons system is defined as in Eq. (7), with the addition of a term $h\left(\xi_{1}^{(1)}\phi_{1}+\xi_{2}^{(1)}\phi_{2}\right)$. We use a gain $\gamma=3.4$, and a field strength $h=0.52$. The resulting equations are integrated using a forward Euler scheme with time step $\Delta t=0.01$ for a total of $T=2000$ steps. The trajectories shown in Fig. 1 (b–d, f–h) start from the same initial condition $\{x_{1}=0.1,\,x_{2}=0.4\}$, which leads to the same overlap against both the ”spurious” and target patterns. In this way, the ”spurious” pattern can significantly interfere with the retrieval process. The short-term plasticity strength is taken to be $k=0$ for Fig. 1 (f–h) and $k=3$, with short-term plastic relaxation time $p=1.0$, for Fig. 1 (b–d). The three panels in each triptych correspond to times $t\simeq 0.01$, $t\simeq 1.0$, and $t\simeq 4.0$.
Appendix B Lyapunov function of RNN with short-term synaptic plasticity
@@
We first differentiate the WW–dependent term. We obtain
@@
Appendix C Cavity Calculation of Global Minima
In this Appendix, we use the static cavity method [39, 9] to obtain a set of self-consistent equations for the fixed point of the recurrent neural network with short-term synaptic plasticity and long-term Hebbian rule. Our starting point are the equation of motion for the action potential of the neurons in the system and the plasticity matrix, given by Eq. (1) and Eq. (5). These equations are recapitulated here for completeness, and are given by
@@ Appendix C Cavity Calculation of Global Minima
C.1 Breakdown of the cavity and spin-glass transition
@@
Expression for the mean squared susceptibilities $\overline{\chi}^{(m,s)}_{2}$, $\overline{\chi}^{(\phi,h)}_{2}$ can be derived through a similar procedure to the one followed in Sec. C for the derivation of the susceptibilities. The final result is
@@
C.2 Phase boundaries with short-term synaptic plasticity
@@
We now address how short-term synaptic plasticity affects the shape of the boundary $\gamma_{g}(\alpha)$ between the paramagnetic phase and the spin glass phase. This transition corresponds to the point where the single-site cavity description breaks down due to the divergence of the fragility $\Xi$, as discussed in Appendix C.1. The condition for this breakdown is given by Eq. (64), when the inequality is saturated. In the paramagnetic phase ($m^{*}=q=0$), the membrane potential of a cavity neuron is $x=\Gamma\phi$ (see Eqs. 9, 10, 10, 11, 12, 13 and 14). Since the membrane potential is independent from the realizations of the Gaussian field $z$, we have that $\left[\overline{\chi}^{(\phi,h)}\right]^{2}=\overline{\chi}_{2}^{(\phi,h)}$. Therefore, the instability condition in Eq. (64) becomes
@@
Appendix D DMFT Derivation
@@
since $A_{ij}(0)=0$, only the non-local self-interaction term survives, similarly to the results obtained by [7]. Inserting Eq. (77) into Eq. (1), we obtain the following expression:
@@
Since we already dealt with the synaptic term, this can be left out in the following computations. We will focus on the term with $J_{ij}$ only, and put back the synaptic kernel in the end. As done in App. C, we introduce a dynamical overlap for each pattern $\mu=1,\ldots,P$,
@@
and we split the set of patterns $\xi_{i}^{\mu}$ into a set of condensed patterns, denoted without loss of generality by $\{\mu^{\ast}\}=\{1,\dots,s\}$, and a set of uncondensed patterns, for $\mu>s$. The overlap against condensed patterns are of order $m^{\mu^{*}}(t)\sim O(1)$, while the overlaps against the uncondensed patterns are of order $m^{\mu}(t)\sim O(N^{-1/2})$. We are going to add in two distinct steps two distinct cavity variables (a neuron and a pattern respectively, as done in App. C) to the dynamical evolution of a neuron in the absence of short-term synaptic plasticity:
@@
D.1 Final self-consistent DMFT with plastic kernel